OpenStudy (anonymous):

I suck with division, can you help me through this problem? x^2-81 / x+9

OpenStudy (anonymous):

factor the top and u will c the answer

OpenStudy (anonymous):

\[\frac{x^2-81}{x+9}=\frac{(x+9)(x-9)}{x+9}=...\]

OpenStudy (anonymous):

how do you factor it?

OpenStudy (anonymous):

hickninja factored it for u

OpenStudy (anonymous):

oh, ok. thank you.

OpenStudy (anonymous):

difference of two perfect squares: \[a^2-b^2=(\sqrt{a} + \sqrt{b})(\sqrt{a}-\sqrt{b})\]

OpenStudy (anonymous):

do the x+9 cancel out? and then the answer is x-9?

OpenStudy (anonymous):

dingdingding!

OpenStudy (anonymous):

ok, thank you for your help I will post another for some help on it.